Shift Summary for owl shift, 24/sep/99

AG (ag@fnal.gov)
Fri, 24 Sep 1999 08:08:15 -0400 (EDT)

Shift Summary for owl shift, 24/sep/99

On Shift: SG, AG

Hours of beam available at all intensities: 7h 25min
Hours of beam available at nominal intensity: 7h 25min
Hours of E832 data (special runs not included): 7h 14min
Hours of special runs (laser scans, muon runs...): 0h 0min

Accounting of lost time:

Laser illumination and pipe-line problems.

========================================================================

Runs ended this shift:

Run 14702 - 5.6M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

Stable beam and smooth e799 data taking.

========================================================================

Description of Shift:

run____spill______time_______comment

---------------------------------------------------------

14702___115_____0235____Laser illumination plot shows large empty spot
____________________________at top-left corner, which stays there for a
____________________________few spills. Trying to reset pipe-line, but
____________________________Pipe-line software doesn't work.
____________________________We clean up pipe-line and
____________________________make drt_fresh_start. Meanwhile beam goes
____________________________away for ~5 minutes.

14703_____1_____0251____Start new run. Laser illumination is back to
____________________________normal.

14703___192_____0709____So far so good, and day shift crew arrives.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

o Continue taking data at 8e12

o After 10:00 go to 10e12 (consult with BobT first)